How much energy is required to heat 20.0 grams of Tin and Calcium from 10 oC to 70 oC?

true or false
1
Expert's answer
2017-11-22T13:09:06-0500
Q=mc∆t
Q(Sn)=0.02kg*230j/kg*C*60C=276J
Q(Ca)=0.02kg*623.9J/kg*C=748.68J
Q=276+748.68=1024.68